# Semantic Exploration and Dense Mapping of Complex Environments using Ground Robots Equipped with LiDAR and Panoramic Camera

### v1

@@ IV-B3 Map Adaptation and Refinement
Errors in the data association process can lead to two common issues: a single object’s measurements being split across different object models, or measurements from two different objects being assigned to the same object model. We propose two strategies to address these issues: model merging and model re-centering. For the former case, model merging is applied when two object models have an Intersection over Minimum (IoMin) greater than a threshold $T_{IOM}$. In this scenario, their data are merged to form a new object model. The probabilities of belonging to the object model are averaged, and the observation sets and key frames from both models are concatenated. To address the latter case, we apply model re-centering. During each model update iteration, both the geometric center and the center of mass are recalculated. If the distance between the two exceeds a threshold $T_{doc}$, the model is re-centered to its center of mass, discarding any information outside the new map boundary. This adjustment helps to ”pull back” the model when incorrect measurements have been associated and have biased the map. After finishing all these steps, the dense point cloud from LiDAR SLAM will be registered and stored in each voxel of a model. Filtering out voxels with too sparse a point cloud, which could be noise, a dense object map is extracted.
@@ IV-C Priority-driven Decoupled Local Sampler
During each planning iteration, only active exploration regions that contain uncovered frontiers are considered. Among these, the local sampler is invoked only for regions that intersect with the robot’s Local Planning Horizon (LPH), selecting suitable viewpoints for exploration, as illustrated in Fig. 3(a) . This section is divided into two parts: the first describes the generation, updating and management of semantic and geometric viewpoints; the second explains how to construct a set of viewpoints that collectively cover the local exploration region and its associated observation spaces.
Figure 3: Overview of Priority-driven Decoupled Local Sampler. (a) Sampling semantic viewpoints and geometric viewpoints in a local active exploration region. (b) Semantic viewpoints sampling along one observation space. The viewpoint with the best observation score ($2^{nd}$ viewpoint from left to right) is selected. Viewpoint with incomplete observation($1^{st}$ viewpoint from left to right) is invalid.
@@ IV-C1 Semantic and Geometric Viewpoint Management
As described in Sections III-C and IV-B, the robot must observe targets from all accessible observation spaces. To achieve this, semantic viewpoints are sampled along the centerline of each observation space at a fixed spatial resolution, as shown in Fig. 3(b). For each center line, the candidate viewpoint with the highest observation score is selected. The score is calculated by projecting the object model into the camera frame and taking the area of the model’s mask. To ensure observation quality and completeness, two types of viewpoints are discarded: (a) those resulting in occluded object observations, and (b) those yielding incomplete object observations, determined via ray-casting from the current pose. In case (b), an additional viewpoint is sampled along the same direction with one higher resolution increment. If case (b) occurs, the sampler will sample an additional viewpoint along the same direction with one more resolution increment. Viewpoints are dynamically updated following two rules: (a) If the robot passes by an observation space and reaches an observation score above a threshold $T_{obs}$ , the corresponding viewpoint is removed; and (b) If a viewpoint falls into a voxel that changes from unknown to occupied, it is discarded and resampled along the centerline.
After updating, the sampled viewpoints are registered in a 2D grid with resolution $g$, with each grid cell assigned to its corresponding active exploration region. The planner considers only the grid centers as viewpoints. This reduces redundant visits to nearby viewpoints—especially those associated with different objects—while maintaining sufficient multi-view coverage.
Geometric viewpoints are obtained using a hierarchical Probabilistic RoadMap (PRM). At the lower level, a dense PRM is constructed to support fine-grained local path planning. At the higher level, nodes are selectively sampled from the dense PRM to serve as geometric viewpoints. Each node is assigned an information gain (IG) score, defined as the number of unknown voxels visible from its pose, given the sensor model. To improve efficiency, IG is approximated using sparse ray-casting.
@@ IV-C2 Semantic-Aware Prioritized Sampling
The sampling of geometric and semantic viewpoints must address two key factors: the potential mutual overlap between viewpoints (known as the submodularity property [7]) and the need for specific semantic viewpoints for object observation. To account for these, we propose a semantic-prioritized sampling mechanism incorporating viewpoints’ mutual overlap into consideration. Our approach begins by adding all frontiers within the current exploring region to a frontier queue. To ensure effective multi-view observation, our sampler gives precedence to semantic viewpoints, pushing them to the viewpoints queue and calculating their frontier coverage. The covered frontiers are then removed from the queue. Subsequently, geometric viewpoints receive NBV-reward values based on expected information gain over time.:
@@
All geometric viewpoints are pushed into an NBV-reward priority queue and are sampled in descending order. Each viewpoint undergoes pseudo-sampling, where the sampler applies its sensor model to estimate frontier coverage. The sampler accepts viewpoints with significant incremental coverage, applying their effects to the frontier queue while dropping those with minimal impact. The process terminates when either the reward queue is empty or the current sample’s reward becomes negligible.
@@ 3       AggressiveGraphConstruction()
5      $success,\sigma_{path}\leftarrow\textbf{exploreReplan()}$()
@@ IV-D Global Planner
The global planner routes through all active regions containing uncovered frontiers, operating in a Receding-Horizon approach—executing only one global step before recalculating. We formulate this as an Asymmetric Traveling Salesman Problem (ATSP) where nodes represent viewpoints sampled by the local sampler. The cost matrix $\mathbf{C}$ has the dimension of $(N+1,N+1)$, where N is the number of viewpoints. Each entry $\mathbf{C_{ij}}$ represents the traveling time cost from Viewpoint $i$ to Viewpoint $j$:
$\mathbf{C_{i,j}}=\mathbf{C_{j,i}}=t^{i,j}_{est},i,j\in\{1,2,\ldots,N\},$ (6)
where $t^{i,j}_{est}$ is the estimated travel time cost from viewpoint $i$ to viewpoint $j$. Unlike traditional approaches, we use different cost metrics for different viewpoints. Since the global path is executed in a Receding-Horizon manner, only the first segment is executed precisely. The travel time from the current position to the $k^{th}$ viewpoint is computed as:
@@
However, since most viewpoints are located in an unknown space, their shortest paths cannot be precisely determined and may never be executed. Therefore, the planner employs Euclidean Distance to evaluate travel cost:
where $VP$ is the viewpoint, c is a time-cost-coefficient. Since the TSP typically involves a closed loop, but exploration requires an open path, the planner creates a virtual loop by assigning zero cost from any viewpoint back to the current position, formulating an ATSP:
$\mathbf{C_{k,0}}=0.$ (9)
After solving the global path, only the segment to the first goal is executed. The planner then replans until all viewpoints in active regions are covered, marking the end of exploration.
IV-E Safe Aggressive Exploration State Machine
Traditional planners restrict goals and paths to known free space for safety, but they lead to short-sighted goal selection. Conversely, naively assuming unknown space is traversable risks instability through overshooting or ”pseudo-collisions” (false collision perceptions causing unexpected halts). To balance speed and safety, we propose an Aggressive Exploration State Machine with three states: $\{exploring,executing,recovering\}$, maintaining planning continuity while allowing rapid exploration.
(1) In the $exploring$ state, the planner aggressively constructs and dynamically maintains a locally safe PRM graph by Aggressive Graph Construction (AGC). It samples nodes in both unknown and free space, assuming that all non-occupied edges are valid. The planner performs A* path search and shortcuts using this same aggressive assumption. We adopt an adaptive PRM sampling strategy following [24] that prioritizes areas with sparse nodes. As the robot moves, the planner dynamically updates the dense PRM by identifying newly occupied nodes, removing them along with their neighbors and edges inside the LPH at a certain frequency. Unlike [14], our approach dynamically updates the PRM to account for mistaken assumptions about unknown space, maintaining a locally safe PRM during aggressive sampling. Upon finding a valid exploration path, the system transitions to the $executing$ state. Otherwise, the system identifies a ”pseudo collision” and enters the $recovering$ state.
(2) During $executing$ state, the system conservatively verifies whether the path and its surrounding areas (within a safe distance) consist of actual free space, checking in backward order. If deemed unsafe, the system reverts to the exploring state to generate a new path.
(3) In the $recovering$ state, we propose a backward recovery mechanism to guide the robot away from the pseudo-collision point and resume exploration. Once the recovery state is activated, the robot backtracks along its last valid exploration path. To ensure that it retreats to a sufficiently distant location, the robot iteratively extends the recovery trajectory backward until it reaches a predefined threshold $T_{\text{back}}$. Simultaneously, the $exploring$ state is reactivated in parallel to proactively plan a new exploration path.
@@ V-A Implementation Details
We configure exploration parameters based on environment scale: 15 m grid size with 8 m LiDAR effective range for large-scale simulations, 5 m grid and 3.5 m range for construction sites, and 10 m grid with 5.0 m range for lobby scenarios. All setups use N=8 viewpoints, 0.1 m resolution for occupancy maps and coarse object models, and FAST LIO2[25] for dense point cloud mapping.
@@
Scene Metric Semantic Eight HIRE Semantic Ours
Office Runtime (s) $536.2\pm 81.1$ $613.7\pm 59.5$ $\mathbf{326.1\pm 11.3}$
Path-length (m) $370.7\pm 46.8$ $380.6\pm 34.6$ $\mathbf{197.6\pm 6.2}$
Map(%) $94.7\pm 0.4$ $\mathbf{96.3\pm 0.8}$ $95.3\pm 1.0$
Viewpoints (%) $99.6\pm 0.8$ $97.3\pm 3.7$ $\mathbf{100.0\pm 0.0}$
Warehouse Runtime (s) $1092.0\pm 72.7$ $1287.0\pm 194.9$ $\mathbf{690.5\pm 30.3}$
Path-length (m) $738.6\pm 58.3$ $1013.2\pm 144.9$ $\mathbf{444.1\pm 20.4}$
Map (%) $93.3\pm 1.8$ $94.2\pm 0.3$ $\mathbf{94.6\pm 0.7}$
Viewpoints (%) $97.1\pm 2.0$ $\mathbf{100.0\pm 0.0}$ $\mathbf{100.0\pm 0.0}$
Factory Runtime (s) $1110.9\pm 119.8$ $1247.1\pm 215.2$ $\mathbf{840.2\pm 32.8}$
Path-length (m) $693.5\pm 108.8$ $858.4\pm 154.0$ $\mathbf{519.5\pm 55.2}$
Map (%) $94.8\pm 0.16$ $\mathbf{96.4\pm 0.8}$ $95.9\pm 0.98$
Viewpoints (%) $97.5\pm 5$ $98.1\pm 2.3$ $\mathbf{100.0\pm 0.0}$
  • 1
    Map(%) means the percentage of map coverage. View(%) is the percentage of observed semantic viewpoints over total number of expected semantic viewpoints.
V-B Benchmark Comparisons
To evaluate the performance of our hierarchical planner, we performed benchmark experiments in three realistic industrial environments in Isaac Sim: 1) Office (40$\times$40$\times$5 m) with 13 chairs as target objects , Warehouse (40$\times$60$\times$5 m) with 9 forklifts as target objects and Factory (45$\times$70$\times$5 m) with 11 robot arms and forklifts as target objects. To demonstrate cross-platform applicability, we used a Husky UGV with a velocity range of 0.5 m/s to 2.2 m/s. The robot is equipped with one LiDAR and two cameras (front and rear) to simulate 360° sensing. To isolate planning performance, we assume ground-truth 3D object detections.
We compared our approach with HIRE, a recent fast exploration algorithm [14], and Semantic Eight [1], a recent semantic exploration algorithm. To ensure fairness and adapt their metrics to align with our task definition, we enhanced HIRE with a semantic weighting scheme where information gain is calculated as the sum of spatial IG and the semantic gain derived from unvisited visible voxels. Since Semantic Eight was originally designed for depth cameras and tightly integrated with a camera-based SLAM system, we re-implemented its planner to suit our LiDAR-based sensor configuration. Each experiment was repeated five times in each environment. As outlined in Table I, our planner consistently outperformed HIRE, achieving faster exploration times, shorter path lengths, and reduced standard deviations, indicating greater robustness. The trajectory visualizations in Fig. 4 confirm that our method shows a much more consistent and efficient trajectory to finish the coverage with less redundancy across all three scenes.
@@ V-C Ablation Study
We conducted an ablation study in three environments to evaluate the effectiveness of our Safe Aggressive Exploration State Machine, consisting of two key components: Aggressive Graph Construction (AGC) and the Safe State Machine (SM). We evaluated how AGC improves exploration efficiency, how naive AGC without SM falls into collision, and how SM ensures the safety of AGC. As shown in Table II, comparing columns 3 and 5 demonstrates that AGC significantly reduces the path length and decreases the time cost in most scenarios, confirming its contribution to improved planner efficiency. In the Factory scene, the AGC+SM strategy produces a slightly longer time but shorter path length. This occurs because Factory is so crowded and complicated that aggressive goal selection triggers lots of stops and recoveries, costing more time than conservative behavior. Furthermore, a comparison between columns 3 and 4 shows that using AGC without SM results in failures during long-term exploration in cluttered environments, underscoring the critical safety role of the SM component.
TABLE II: Ablation study of Safe Aggressive Exploration State Machine
Scene Metric AGC+SM Only AGC Only SM
Office Time (s) $\mathbf{326.1\pm 11.3}$ 175.8 $433.7\pm 26.9$
Path (m) $\mathbf{197.6\pm 6.2}$ 89.1 $270.7\pm 30.0$
Map (%) $95.3\pm 1.0$ 52.8 $\mathbf{96.1\pm 0.3}$
View (%) $\mathbf{100.0\pm 0.0}$ 50.8 $\mathbf{100.0\pm 0.0}$
Warehouse Time (s) $\mathbf{690.5\pm 30.3}$ 343.1 $742.2\pm 77.9$
Path (m) $\mathbf{444.1\pm 20.3}$ 219.2 $469.2\pm 53.7.4$
Map (%) $\mathbf{94.6\pm 0.7}$ 66.5 $93.6\pm 1.2$
View (%) $\mathbf{100.0\pm 0.0}$ 55.1 $\mathbf{100.0\pm 0.0}$
Factory Time (s) $840.2\pm 32.8$ 327.2 $\mathbf{813.5\pm 28.5}$
Path (m) $\mathbf{519.5\pm 55.2}$ 205.2 $542.8\pm 43.9$
Map (%) $95.9\pm 0.98$ 52.3 $\mathbf{96.4\pm 1.3}$
View (%) $\mathbf{100.0\pm 0.0}$ 34.4 $\mathbf{100.0\pm 0.0}$
  • 1
    is the mean value from incomplete runs; the system did not successfully complete in any trial.

### v2

@@ IV-B3 Map Adaptation and Refinement
Errors in the data association process can lead to two common issues: a single object’s measurements being split across different object models, or measurements from two different objects being assigned to the same object model. We propose two strategies to address these issues: model merging and model re-centering. For the former case, model merging is applied when two object models have an Intersection over Minimum (IoMin) greater than a threshold $T_{IOM}$. In this scenario, their data is merged to form a new object model. The probabilities of belonging to the object model are averaged, and the observation sets and key frames from both models are concatenated. To address the latter case, we apply model re-centering. During each model update iteration, both the center of geometry and mass are recalculated. If their distance exceeds a threshold $T_{doc}$, the model is re-centered to its center of mass, discarding any information outside the new map boundary. This adjustment helps to ”pull back” the model when incorrect measurements have been associated and have biased the map. Finally, the dense point cloud from LiDAR SLAM will be registered and stored in each voxel of a model. Filtering out voxels with too sparse a point cloud, a dense object map is extracted.
@@ IV-C Priority-driven Decoupled Local Sampler
During each planning iteration, only active exploration regions that contain uncovered frontiers are considered. Among these, the local sampler is invoked only for regions that intersect with the robot’s Local Planning Horizon (LPH), selecting suitable viewpoints for exploration, as illustrated in Fig. 3(a) . This section is divided into two parts: the first describes the generation, updating, and management of semantic and geometric viewpoints; the second explains how to construct a set of viewpoints that collectively cover the local exploration region and associated observation spaces.
Figure 3: Overview of Priority-driven Decoupled Local Sampler. (a) Sampling semantic and geometric viewpoints in a local active exploration region. (b) Semantic viewpoint sampling along one observation space. The viewpoint with the best observation score ($2^{nd}$ viewpoint from left to right) is selected. Viewpoint with incomplete observation ($1^{st}$ viewpoint) is discarded.
@@ IV-C1 Semantic and Geometric Viewpoint Management
As described in Sections III-C and IV-B, the robot must observe targets from all accessible observation spaces. To achieve this, semantic viewpoints are sampled along the centerline of each observation space at a fixed spatial resolution, as shown in Fig. 3(b). For each center line, the candidate viewpoint with the highest observation score is selected. The score is calculated by projecting the object model into the camera frame and taking the area of the model’s mask. To ensure observation quality and completeness, two types of viewpoints are discarded: (a) those resulting in occluded object observations, and (b) those yielding incomplete object observations, determined via ray-casting from the current pose. Fig. 3(b) shows an example of viewpoint type (b), where the first viewpoint from left to right will produce an incomplete observation. The sampler will sample further viewpoints along the same direction with a fixed resolution interval and pick the one with the best observation score. Viewpoints are dynamically updated following two rules: (a) if the robot passes by an observation space and reaches an observation score above a threshold $T_{obs}$, the corresponding viewpoint is removed; and (b) if a viewpoint falls into a voxel being updated to occupied, it is discarded, and an alternative viewpoint will be selected by repeating the above process. After updating, the viewpoints are registered in a 2D grid with resolution $g$ and assigned to the corresponding active exploration region. The planner considers only the grid centers as viewpoints to reduce redundant visits to nearby viewpoints when the distribution is crowded.
Geometric viewpoints are obtained using a hierarchical Probabilistic RoadMap (PRM). At the lower level, a dense PRM is constructed to support fine-grained local path planning. At the higher level, nodes are selectively sub-sampled from the dense PRM by maintaining a controlled node density to serve as geometric viewpoints. Only the sampled viewpoints are assigned an information gain (IG) score, defined as the number of unknown voxels visible from its pose.
@@ IV-C2 Semantic-Aware Prioritized Sampling
The sampling of geometric and semantic viewpoints must address two key factors: the potential mutual overlap between viewpoints (known as the submodularity property [8]) and the need for specific semantic viewpoints for object observation. To account for these, we propose a semantic-prioritized sampling mechanism incorporating viewpoints’ mutual overlap into consideration. Our approach begins by adding all frontiers within the current exploring region to a frontier queue. To ensure effective multi-view observation, our sampler gives precedence to semantic viewpoints, pushing them to the viewpoints queue and calculating their frontier coverage. The covered frontiers are then removed from the queue. Subsequently, geometric viewpoints receive NBV-reward based on accumulated information gain:
@@
All geometric viewpoints are pushed into an NBV-reward priority queue and are sampled in descending order. Each viewpoint undergoes pseudo-sampling to apply its corresponding frontier coverage. The sampler accepts viewpoints with significant incremental coverage, applying their effects to the frontier queue while dropping those with minimal impact. The process terminates when either the reward queue is empty or the current sample’s reward becomes negligible.
@@ 3 AggressiveGraphConstruction()
5 $success,\sigma_{path}\leftarrow\textbf{exploreReplan()}$
@@ IV-D Global Planner
The global planner routes through all active regions containing uncovered frontiers, operating in a receding-horizon approach—executing only one global step before recalculating. We formulate this as an Asymmetric Traveling Salesman Problem (ATSP) similar as [7]. Unlike [7], since only the first path segment is executed precisely, the travel time from the current position to the $k^{th}$ viewpoint is computed as:
@@
where $n$ is the number of nodes in the shortest path found by the $A^{*}$ algorithm. The travel costs among other viewpoints, which may even be located in unknown spaces, are evaluated by Euclidean Distance:
where $VP$ is the viewpoint, and c is a time-cost coefficient.
The planner keeps recomputing in a receding-horizon manner until all viewpoints in active exploration regions are covered, marking the end of exploration.
IV-E Safe Aggressive Exploration State Machine
@@
(1) In the exploringexploring state, the planner aggressively constructs and dynamically maintains a locally safe PRM graph by Aggressive Graph Construction (AGC). It samples nodes in both unknown and free space, assuming all non-occupied connecting edges are valid. The planner performs A* path search and shortcuts using the same assumption. We adopt an adaptive PRM sampling strategy following [29] that prioritizes areas with fewer nodes. As the robot moves, the planner dynamically updates the dense PRM by removing newly occupied nodes with associated edges inside the LPH at a certain frequency. Unlike [14], our approach dynamically updates the PRM to account for mistaken assumptions about unknown space, maintaining a locally safe PRM during aggressive sampling. Upon finding a valid exploration path, the system transitions to the executingexecuting state. If no path exists, the system identifies a ”pseudo collision” and enters the recoveringrecovering state.
(2) During executingexecuting state, the system conservatively verifies whether the path and its surrounding areas (within a safe distance) consist of actual free space, checking in backward order. If deemed unsafe, the system reverts to the exploring state to generate a new path.
(3) In the recoveringrecovering state, we propose a backward recovery mechanism to guide the robot away from the pseudo-collision point and resume exploration. Once the recovery state is activated, the robot backtracks along its last valid exploration path. To ensure that it retreats to a sufficiently distant location, the robot iteratively extends the recovery trajectory backward until it reaches a predefined threshold $T_{\text{back}}$. Simultaneously, the exploringexploring state is reactivated in parallel to proactively plan a new exploration path.
@@ V-A Implementation Details
We configure exploration parameters based on environment scale: 15 m grid size with 8 m LiDAR effective range for large-scale simulations, 5 m grid and 3.5 m range for construction sites, and 10 m grid with 5.0 m range for lobby scenarios. We use a safe distance of 0.1 m as path planning shortcut tolerance. All setups use N=8 viewpoints, 0.1 m resolution for occupancy maps and coarse object models, and FAST LIO2 [30] for real-time background SLAM.
V-B Benchmark Comparisons
[…]
@@ V-B Benchmark Comparisons
Scene Method Runtime (s) Path Length (m) Map Comp.a (%) VP Comp.b (%)
Office SE [1] $536.2\pm 8.1$ $370.7\pm 4.6$ $94.7$ $99.6$
HIRE [14] $613.7\pm 59.5$ $380.6\pm 34.6$ $\mathbf{96.3}$ $97.3$
TARE [8] $598.0\pm 55.7$ $340.6\pm 34.4$ $94.8$ $66.0$
Ours $\mathbf{326.1\pm 11.3}$ $\mathbf{197.6\pm 6.2}$ $95.3$ $\mathbf{100.0}$
Warehouse SE [1] $1092.0\pm 72.7$ $738.6\pm 58.3$ $93.3$ $97.1$
HIRE [14] $1287.0\pm 194.9$ $1013.2\pm 144.9$ $94.2$ $\mathbf{100.0}$
TARE [8] $863.0\pm 48.6$ $485.1\pm 23.6$ $92.1$ $82.5$
Ours $\mathbf{690.5\pm 30.3}$ $\mathbf{444.1\pm 20.3}$ $\mathbf{94.6}$ $\mathbf{100.0}$
Factory SE [1] $1110.9\pm 119.8$ $693.5\pm 108.8$ $94.8$ $97.5$
HIRE [14] $1247.1\pm 215.2$ $858.4\pm 154.0$ $\mathbf{96.4}$ $98.1$
TARE [8] $900.6\pm 45.8$ $530.8\pm 30.5$ $90.8$ $76.4$
Ours $\mathbf{840.2\pm 32.8}$ $\mathbf{519.5\pm 55.2}$ $95.9$ $\mathbf{100.0}$
  • a
    Percentage of voxel map coverage.
  • b
    Percentage of observed semantic viewpoints over total expected number.
We compared our approach with HIRE, a recent fast exploration algorithm [14], and Semantic Eight (SE) [1], a recent semantic exploration algorithm [^1] Since Semantic Eight was originally designed for depth cameras and tightly integrated with a camera-based SLAM system, we re-implemented its planner to leverage our LiDAR-based sensor configuration.. To ensure fairness and adapt their metrics to align with our task definition, we enhanced HIRE with a semantic weighting scheme where information gain is calculated as the weighted sum of spatial IG and the semantic gain derived from unvisited visible voxels. In addition, we choose TARE [8], a well-known ground robot exploration algorithm, as a baseline to validate our performance in geometric exploration. Each experiment was repeated five times in each environment. As detailed in Table I, our planner consistently outperformed all of these methods, achieving faster exploration times, shorter path lengths, and reduced standard deviations, indicating greater robustness. The trajectory visualizations in Fig. 4 confirm that our method shows a much more consistent and efficient trajectory to finish the coverage across all three scenes.
@@ V-C Ablation Study
We conducted an ablation study in three environments to evaluate the effectiveness of our Priority-driven Decoupled Local Sampler (PDLS) and Safe Aggressive Exploration State Machine (SAESM).
To validate PDLS, we conducted comparisons against a local greedy-information-gain sampler (LGS). It samples the viewpoint with the largest gain greedily in the local active exploration region while keeping all other modules the same as ours222 LGS strategy generates longer paths and the robot tends to move at a higher speed compared with PDLS. To ensure the completion of the exploration task, we use a safe distance of 0.3 m for local-greedy sampling.. To incorporate the semantic coverage, we define the overall gain as the weighted sum of the geometric gain and semantic gain derived from unvisited visible voxels of semantic objects. Comparing the Proposed (PDLS) with LGS in Table. II, we achieve higher semantic and geometric coverage rates.
The SAESM consists of two key components: Aggressive Graph Construction (AGC) and the Safe State Machine (SM). We evaluated the effectiveness of each. As shown in Table II, comparing No-AGC and Proposed demonstrates that AGC significantly reduces the exploration time in most scenarios, confirming its contribution to improved planner efficiency. In the Factory scene, the proposed strategy produces a slightly longer time but shorter path length. This occurs because Factory is so crowded and complicated that aggressive goal selection triggers a lot of stops and recoveries, costing more time than conservative behavior. Furthermore, a comparison between No-SM and Proposed shows that using AGC without SM results in task failures during cluttered environments, underscoring the critical safety role of the SM component.
TABLE II: Ablation study
Scene Method Runtime (s) Path Length (m) Map Comp. (%) VP Comp. (%)a
Office No-SM $175.8^{\dagger}$ $89.1^{\dagger}$ $52.8^{\dagger}$ $50.8^{\dagger}$
No-AGC $433.7\pm 26.9$ $270.7\pm 30.0$ $\mathbf{96.1}$ $\mathbf{100.0}$
LGS $496.4\pm 78.7$ $272.5\pm 26.1$ $94.0$ $\mathbf{100.0}$
Proposed $\mathbf{326.1\pm 11.3}$ $\mathbf{197.6\pm 6.2}$ $95.3$ $\mathbf{100.0}$
Warehouse No-SM $343.1^{\dagger}$ $219.2^{\dagger}$ $66.5^{\dagger}$ $55.1^{\dagger}$
No-AGC $742.2\pm 77.9$ $469.2\pm 53.7$ $93.6$ $\mathbf{100.0}$
LGS $874.6\pm 205.0$ $503.6\pm 47.7$ $91.6$ $99.4$
Proposed $\mathbf{690.5\pm 30.3}$ $\mathbf{444.1\pm 20.3}$ $\mathbf{94.6}$ $\mathbf{100.0}$
Factory No-SM $327.2^{\dagger}$ $205.2^{\dagger}$ $52.3^{\dagger}$ $34.4^{\dagger}$
No-AGC $840.2\pm 32.8$ $542.8\pm 43.9$ $\mathbf{96.4}$ $\mathbf{100.0}$
LGS $946.3\pm 125.8$ $585.8\pm 56.5$ $93.2$ $99.3$
Proposed $\mathbf{813.5\pm 28.5}$ $\mathbf{519.5\pm 55.2}$ $95.9$ $\mathbf{100.0}$
  • $\dagger$
    presents the data from incomplete runs; the system failed in all trials.